# Robust phase estimation of the ground-state energy without controlled time evolution on a quantum device

### v1

@@ Abstract
Introduction— In quantum many-body systems and quantum chemistry, determining the ground-state energy is often important. However, since the Hilbert space dimension of the Hamiltonian grows exponentially with the number of constituent elements (eg, the number of electrons), it becomes challenging to compute these eigenvalues by numerically diagonalizing the Hamiltonian using classical computers. To address this issue, several algorithms have been proposed to estimate energy eigenvalues of quantum many-body Hamiltonians using quantum computers, including those designed for noisy intermediate-scale quantum devices [1, 2, 3, 4, 5, 6, 7, 8, 9], early fault-tolerant quantum computer (FTQC) [10, 11, 12, 13, 14], and fully FTQC [15, 16, 17, 18]. One approach to estimate the eigenvalues of the Hamiltonian $\hat{H}$ is the phase estimation algorithm. In this method, an initial state is prepared by using adiabatic state preparation (ASP). The eigenvalues of the Hamiltonian are encoded in the phases of quantum states by applying a controlled version of the time evolution operator $\hat{U}=e^{-i\hat{H}t}$, which causes the quantum state to accumulate a relative phase that corresponds to the energy eigenvalue of the Hamiltonian. By measuring the phases of the quantum states after evolution using the phase estimation algorithm, one can extract the eigenvalues of the Hamiltonian [19, 20, 21, 22, 23, 24, 25]. However, combining ASP with phase estimation involves controlled time-evolution operations, which can lead to prohibitively high implementation costs. To overcome this challenge, subsequent research proposed methods for estimating energy gaps without controlled time evolution [26, 27, 28, 29, 30]. In particular, methods proposed in Refs. [29, 30] use ASP and Ramsey-type measurements to estimate the energy differences. Although such methods can estimate the energy differences of the eigenvalues, directly estimating the eigenvalues remains challenging.
@@
Method—
Figure 1: (a) A conceptual diagram of our method. The initial state is a superposition of the ground state of the driving Hamiltonian and a reference state of $\hat{H}_{\mathrm{D}}$. The system’s Hamiltonian $\hat{H}(t)$ is expressed as a time-dependent combination of the driving and problem Hamiltonians, as shown in Eq (1), with the function $F(t)$ expressed in Eq. (2). For $0\leq t\leq T$, the Hamiltonian corresponds to the standard ASP Hamiltonian, denoted as $\hat{H}_{\mathrm{ASP}}$. Subsequently, for $T\leq t\leq T+\tau$, the Hamiltonian is represented as $\hat{H}_{\mathrm{TE}}$, during which time evolution encodes the relative phase corresponding to the energy difference between the ground and the reference states of $\hat{H}_{\mathrm{P}}$ into the superposition state. Finally, for $T+\tau\leq t\leq 2T+\tau$, the Hamiltonian is denoted as $\hat{H}_{\mathrm{RASP}}$, which implements the RASP process. (b) The amplitude $|f(\omega/J)|$ obtained from the numerical simulation as a function of the frequency $\omega/J$. The peaks were observed at $\omega/J\simeq 10.8,8.1,3.8$, and $2.5$. These peaks correspond to the energy differences between the reference state of $\hat{H}_{\mathrm{P}}$ and the ground state , the second excited state, the eighth excited state, and the eleventh excited state, respectively. (c) Comparison of ground-state energy estimation results obtained via conventional ASP and those obtained with our proposed method. The horizontal axis represents the total runtime $2T+\tau$, which is common to both methods, and the vertical axis indicates the estimated ground-state energy. The blue markers represent the ground-state energy estimates obtained using the conventional method, while the red markers show the estimates calculated with our proposed method. The black dashed line represents exact diagonalization result.
@@
For $0\leq t\leq T$, the Hamiltonian is the standard ASP Hamiltonian $\hat{H}_{\mathrm{ASP}}$. During the next interval, $T\leq t\leq T+\tau$, the state evolves under the problem Hamiltonian $\hat{H}_{\mathrm{P}}$, and the superposition of the ground state and the reference state of $\hat{H}_{\mathrm{P}}$ acquires a relative phase corresponding to the energy gap between the two states, which is similar to the Ramsey-type measurement. Finally, during $T+\tau\leq t\leq 2T+\tau$, the system undergoes RASP under the Hamiltonian $\hat{H}_{\mathrm{RASP}}$ [31, 32, 33, 34]. In our method, we make the following assumptions:
@@
For example, quantities such as the total magnetization $\hat{M}=\sum_{i=1}^{N}\hat{\sigma}^{z}_{i}$ can be used as $\hat{Q}$, where $N$ is the number of qubits and $\hat{\sigma}^{z}_{i}$ denotes the $z$-component of the Pauli operator for the $i$-th qubit. When these assumptions are satisfied, $\hat{H}_{\mathrm{D}}$ and $\hat{H}_{\mathrm{P}}$ are block-diagonalized for each value of $q$. For a given eigenvalue $q$ as stated above, the corresponding subspace is of size $O(\mathrm{poly}(N))$, making it diagonalizable using a classical computer. The eigenstates obtained in each subspace are denoted as |$\ket*{\psi_{q,\mathrm{D}}}_{n}$ and |$\ket*{\psi_{q,\mathrm{P}}}_{n}$, where $n$ represents the energy level within the subspace associated with eigenvalue $q$. In our proposed method, we utilize these eigenstates, which we call reference states. We assume that the energy of the reference state |$\ket*{\psi_{q,\mathrm{P}}}_{n}$ has the largest energy in the problem Hamiltonian. If the given problem Hamiltonian does not naturally satisfy this assumption, we can add an additional term such as $H_{\rm{add}}=-\lambda(\hat{Q}-q)^{2}$ where $\lambda$ is a positive hyperparameter. It is worth noting we know the eigenvalues of these reference states using the classical computer. The use of a trivial eigenstate and its superposition with the ground state for estimating the ground-state energy of the quantum chemistry Hamiltonian via ASP was briefly mentioned in [35] but no detailed explanation was provided. On the other hand, our proposed method provides a much more general way to create and use the reference state for the ground-state energy estimation without control time evolution. Using a superposition of the reference state of $\hat{H}_{\mathrm{D}}$ in the subspace of a given eigenvalue $q$ and the ground state of $\hat{H}_{\mathrm{D}}$ in the subspace of a given eigenvalue $q^{\prime}$ as the initial state, we have |$\ket*{\psi_{\mathrm{ini}}}=\frac{1}{\sqrt{2}}(\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}+\ket*{\psi_{q,\mathrm{D}}}_{n})$, where |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}$ is the ground state of the driving Hamiltonian in the subspace of a given eigenvalue $q^{\prime}$. After ASP under $\hat{H}(t)$, the system evolves into: |$\ket*{\psi}=\frac{1}{\sqrt{2}}(\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}+e^{i\theta}\ket*{\psi_{q,\mathrm{P}}}_{n})$, where |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}$ is the ground state of the problem Hamiltonian in the subspace of a given eigenvalue $q^{\prime}$, and $\theta$ is the relative phase arising from the ASP process. Here, |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}$ is the state obtained by performing ASP, which is associated with |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}$. Then, the state evolves under $\hat{H}_{\mathrm{P}}$ for a time $\tau$, leading to $e^{-i\hat{H}_{\mathrm{P}}\tau}\ket*{\psi}=\frac{1}{\sqrt{2}}(e^{-iE^{\mathrm{g}}_{q^{\prime},\mathrm{P}}\tau}\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}+e^{-iE^{n}_{q,\mathrm{P}}\tau+i\theta}\ket*{\psi_{q,\mathrm{P}}}_{n})$, where $E^{n}_{q,\mathrm{P}}$ represents the eigenvalue of |$\ket*{\psi_{q,\mathrm{P}}}_{n}$, which we can obtain using a classical calculator. Subsequently, RASP is executed under $\hat{H}_{\mathrm{RASP}}$, resulting in the final state: |$\ket*{\psi_{\mathrm{f}}}=\frac{1}{\sqrt{2}}(e^{-iE^{\mathrm{g}}_{q^{\prime},\mathrm{P}}\tau}\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}+e^{-iE^{n}_{q,\mathrm{P}}\tau+i\theta^{\prime}}\ket*{\psi_{q,\mathrm{P}}}_{n})$, where $\theta^{\prime}$ is a relative phase, and $E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}$ is the ground-state energy of $\hat{H}_{\mathrm{P}}$. The above sequence is summarized in Fig. 1 (a). A measurement is performed using the projection operator ^=|$\hat{\mathcal{M}}=\outerproduct{\psi_{\mathrm{ini}}}{\psi_{\mathrm{ini}}}$ |, and the probability of projecting onto this state is given by: $\mathcal{P}(\tau)\equiv\absolutevalue{\innerproduct{\psi_{\mathrm{f}}}{\psi_{\mathrm{ini}}}}^{2}=\cos^{2}\quantity[\frac{(E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}})\tau}{2}+\frac{\theta^{\prime}}{2}]$. This signal oscillates with the frequency corresponding to the energy difference. By repeating the above steps and sweeping $\tau$, multiple measurement results are obtained. Finally, a discrete Fourier transform of $\mathcal{P}(\tau)$ is performed as $f(\omega)=\sum_{n=1}^{L}\mathcal{P}(\tau_{n})e^{-i\omega\tau_{n}}$, where the discretized time step $\tau_{n}$ is defined as $\tau_{n}=\tau_{\mathrm{min}}+\frac{n-1}{L-1}(\tau_{\mathrm{max}}-\tau_{\mathrm{min}})$, with $\tau_{\mathrm{min}}$ and $\tau_{\mathrm{max}}$ denoting the minimum and maximum times, and $L$ is the number of steps. However, due to the time-energy uncertainty relation, as the energy gap increases, the time step $\tau_{n}$ must be made smaller. The peak of $f(\omega)$ appears at $\omega$ corresponding to the energy difference $E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}}$. Since $E^{n}_{q,\mathrm{P}}$ is known, the ground-state energy in the subspace corresponding to eigenvalue $q^{\prime}$ can be determined. From this procedure, we successfully estimate the ground-state energy in the sector with $q^{\prime}$. To estimate the ground-state energy in another sector with $q^{\prime\prime}$, we could perform the above procedure, say $q^{\prime\prime}$. Thanks to the second assumption of our protocol, we can perform this search across all sectors within polynomial time. Similarly, the first-excited-state energy in the subspace of $q^{\prime}$ can be estimated by preparing a superposition of the first excited state and the reference state of $\hat{H}_{\mathrm{D}}$ in the subspace of eigenvalues $q^{\prime}$ and $q$ respectively, as the initial state.
@@
Furthermore, if non-adiabatic transitions occur during the ASP, peaks appear in $f(\omega)$ at frequencies corresponding to the energy difference between the excited state and the reference state of $\hat{H}_{\mathrm{P}}$, such as $E^{\mathrm{e_{1}}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}},E^{\mathrm{e_{2}}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}},\dots$, where $E^{\mathrm{e}_{i}}_{q^{\prime},\mathrm{P}}$ denotes the $i$-th excited-state energy in the subspace associated with eigenvalue $q^{\prime}$. These excitation energies can also be estimated since the trivial eigenvalue is known in advance. However, if non-adiabatic transitions significantly reduce the population of the ground state, both our method and the conventional quantum phase estimation fail to determine the ground-state energy. Regarding the reference state, since it is chosen within a subspace of size $O(\mathrm{poly}(N))$, non-adiabatic transitions can be suppressed within polynomial time of $T$. Therefore, non-adiabatic transitions involving the reference state are neglected.
@@
Result— We estimate the ground-state energy of the spin1/212-1/2- 1 / 2 Heisenberg model using our proposed method through numerical simulations. In this study, we consider a 4-qubit system. The problem Hamiltonian is defined as the Heisenberg model with an additional inhomogeneous longitudinal magnetic field to lift the degeneracy:

### v2

@@ Abstract
I Introduction
In quantum many-body systems and quantum chemistry, determining the ground-state energy is often important. However, since the Hilbert space dimension of the Hamiltonian grows exponentially with the number of constituent elements (eg, the number of electrons), it becomes challenging to compute these eigenvalues by numerically diagonalizing the Hamiltonian using classical computers. To address this issue, several algorithms have been proposed to estimate energy eigenvalues of quantum many-body Hamiltonians using quantum computers, including those designed for noisy intermediate-scale quantum devices [1, 2, 3, 4, 5, 6, 7, 8, 9], early fault-tolerant quantum computer (FTQC) [10, 11, 12, 13, 14], and fully FTQC [15, 16, 17, 18]. One approach to estimate the eigenvalues of the Hamiltonian $\hat{H}$ is the phase estimation algorithm. In this method, an initial state is prepared by using adiabatic state preparation (ASP). The eigenvalues of the Hamiltonian are encoded in the phases of quantum states by applying a controlled version of the time evolution operator $\hat{U}=e^{-i\hat{H}t}$, which causes the quantum state to accumulate a relative phase that corresponds to the energy eigenvalue of the Hamiltonian. By measuring the phases of the quantum states after evolution using the phase estimation algorithm, one can extract the eigenvalues of the Hamiltonian [19, 20, 21, 22, 23, 24, 25]. However, combining ASP with phase estimation involves controlled time-evolution operations, which can lead to prohibitively high implementation costs. To overcome this challenge, subsequent research proposed methods for estimating energy gaps without controlled time evolution [26, 27, 28, 29, 30]. In particular, methods proposed in Refs. [29, 30] use ASP and Ramsey-type measurements to estimate the energy differences. Although such methods can estimate the energy differences of the eigenvalues, directly estimating the eigenvalues remains challenging.
@@ I Introduction
II Method
Figure 1: (a) A conceptual diagram of our method. The initial state is a superposition of the ground state of the driving Hamiltonian and a reference state of $\hat{H}_{\mathrm{D}}$. The system’s Hamiltonian $\hat{H}(t)$ is expressed as a time-dependent combination of the driving and problem Hamiltonians, as shown in Eq (1), with the function $F(t)$ expressed in Eq. (2). For $0\leq t\leq T$, the Hamiltonian corresponds to the standard ASP Hamiltonian, denoted as $\hat{H}_{\mathrm{ASP}}$. Subsequently, for $T\leq t\leq T+\tau$, the Hamiltonian is represented as $\hat{H}_{\mathrm{TE}}$, during which time evolution encodes the relative phase corresponding to the energy difference between the ground and the reference states of $\hat{H}_{\mathrm{P}}$ into the superposition state. Finally, for $T+\tau\leq t\leq 2T+\tau$, the Hamiltonian is denoted as $\hat{H}_{\mathrm{RASP}}$, which implements the RASP process. (b) The amplitude $|f(\omega/J)|$ obtained from the numerical simulation as a function of the frequency $\omega/J$. The peaks were observed at $\omega/J\simeq 10.8,8.1,5.4,3.8$, and $2.5$. These peaks correspond to the energy differences between the reference state of $\hat{H}_{\mathrm{P}}$ and the ground state, as well as the second, fifth, seventh, and tenth excited states, respectively. (c) Comparison of ground-state energy estimates obtained using conventional ASP and our proposed method. The horizontal axis represents the total runtime $2T+\tau$ ; for the conventional ASP, we set $T_{\mathrm{conv}}=2T+\tau$ to allow fair comparison. The vertical axis indicates the estimated ground-state energy. The blue markers represent the ground-state energy estimates obtained using the conventional method, while the red markers show the estimates calculated with our proposed method. The black dashed line represents exact diagonalization result.
@@ II Method
For $0\leq t\leq T$, the Hamiltonian is the standard ASP Hamiltonian $\hat{H}_{\mathrm{ASP}}$. During the next interval, $T\leq t\leq T+\tau$, the state evolves under the problem Hamiltonian $\hat{H}_{\mathrm{P}}$, and the superposition of the ground state and the reference state of $\hat{H}_{\mathrm{P}}$ acquires a relative phase corresponding to the energy gap between the two states, which is similar to the Ramsey-type measurement. Finally, during $T+\tau\leq t\leq 2T+\tau$, the system undergoes RASP under the Hamiltonian $\hat{H}_{\mathrm{RASP}}$ [31, 32, 33, 34].
In our method, we make the following assumptions:
@@
For example, quantities such as the total magnetization $\hat{M}=\sum_{i=1}^{N}\hat{\sigma}^{z}_{i}$ can be used as $\hat{Q}$, where $N$ is the number of qubits and $\hat{\sigma}^{z}_{i}$ denotes the $z$-component of the Pauli operator for the $i$-th qubit. When these assumptions are satisfied, $\hat{H}_{\mathrm{D}}$ and $\hat{H}_{\mathrm{P}}$ are block-diagonalized for each value of $q$. It is important to note that one of the subspaces defined by the eigenstates of $\hat{Q}$ contains the global ground state of the problem Hamiltonian. For a given eigenvalue $q$ as stated above, the corresponding subspace is of size $O(\mathrm{poly}(N))$, making it diagonalizable using a classical computer. The eigenstates obtained in that subspace are denoted as |$\ket*{\psi_{q,\mathrm{D}}}_{n}$ and |$\ket*{\psi_{q,\mathrm{P}}}_{n}$, where $n$ represents the energy level within the subspace associated with the eigenvalue $q$. We refer to these eigenstates as “reference states” in our proposed method. Note that our approach does not require all subspaces to be of size $O(\mathrm{poly}(N))$); we only require this property for those subspaces from which we choose our reference states. We assume that the energy of the reference state |$\ket*{\psi_{q,\mathrm{P}}}_{n}$ has the largest energy in the problem Hamiltonian. If the given problem Hamiltonian does not naturally satisfy this assumption, we can add an additional term such as $H_{\rm{add}}=-\lambda(\hat{Q}-q)^{2}$ where $\lambda$ is a positive hyperparameter. It is worth noting that the eigenvalues of these reference states can be determined using a classical computer.
The use of a trivial eigenstate and its superposition with the ground state for estimating the ground-state energy of the quantum chemistry Hamiltonian via ASP was briefly mentioned in [35] but no detailed explanation was provided. On the other hand, our proposed method provides a much more general way to create and use the reference state for the ground-state energy estimation without control time evolution.
Using a superposition of the reference state of $\hat{H}_{\mathrm{D}}$ in the subspace of a given eigenvalue $q$ and the ground state of $\hat{H}_{\mathrm{D}}$ in the subspace of a given eigenvalue $q^{\prime}$ as the initial state, we have
|$\ket*{\psi_{\mathrm{ini}}}=\frac{1}{\sqrt{2}}(\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}+\ket*{\psi_{q,\mathrm{D}}}_{n}),$ (6)
where |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}$ is the ground state of the driving Hamiltonian in the subspace of a given eigenvalue $q^{\prime}$. After ASP under $\hat{H}(t)$, the system evolves into:
|$\ket*{\psi}=\frac{1}{\sqrt{2}}(\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}+e^{i\theta}\ket*{\psi_{q,\mathrm{P}}}_{n}),$ (7)
where |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}$ is the ground state of the problem Hamiltonian in the subspace of a given eigenvalue $q^{\prime}$, and $\theta$ is the relative phase arising from the ASP process. Here, |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}$ is the state obtained by performing ASP, which is associated with |$\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}$. Then, the state evolves under $\hat{H}_{\mathrm{P}}$ for a time $\tau$, leading to
$e^{-i\hat{H}_{\mathrm{P}}\tau}\ket*{\psi}=\frac{1}{\sqrt{2}}(e^{-iE^{\mathrm{g}}_{q^{\prime},\mathrm{P}}\tau}\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{P}}}+e^{-iE^{n}_{q,\mathrm{P}}\tau+i\theta}\ket*{\psi_{q,\mathrm{P}}}_{n}),$ (8)
where $E^{n}_{q,\mathrm{P}}$ represents the eigenvalue of |$\ket*{\psi_{q,\mathrm{P}}}_{n}$, which we can obtain using a classical calculator. Subsequently, RASP is executed under $\hat{H}_{\mathrm{RASP}}$, resulting in the final state:
|$\ket*{\psi_{\mathrm{f}}}=\frac{1}{\sqrt{2}}(e^{-iE^{\mathrm{g}}_{q^{\prime},\mathrm{P}}\tau}\ket*{\psi^{\mathrm{g}}_{q^{\prime},\mathrm{D}}}+e^{-iE^{n}_{q,\mathrm{P}}\tau+i\theta^{\prime}}\ket*{\psi_{q,\mathrm{P}}}_{n}),$ (9)
where $\theta^{\prime}$ is a relative phase, and $E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}$ is the ground-state energy of $\hat{H}_{\mathrm{P}}$. Note that any relative phase acquired during the ASP process does not affect the final result of the eigenvalue estimation (see Appendix A for details). The above sequence is summarized in Fig. 1 (a). A measurement is performed using the projection operator ^=|$\hat{\mathcal{M}}=\outerproduct{\psi_{\mathrm{ini}}}{\psi_{\mathrm{ini}}}$ |, and the probability of projecting onto this state is given by:
$\mathcal{P}(\tau)\equiv\absolutevalue{\innerproduct{\psi_{\mathrm{f}}}{\psi_{\mathrm{ini}}}}^{2}=\cos^{2}\quantity[\frac{(E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}})\tau}{2}+\frac{\theta^{\prime}}{2}].$ (10)
This signal oscillates with the frequency corresponding to the energy difference. By repeating the above steps and sweeping $\tau$, multiple measurement results are obtained. Finally, a discrete Fourier transform of $\mathcal{P}(\tau)$ is performed as
$f(\omega)=\sum_{n=1}^{L}\mathcal{P}(\tau_{n})e^{-i\omega\tau_{n}},$ (11)
where the discretized time step $\tau_{n}$ is defined as $\tau_{n}=\tau_{\mathrm{min}}+\frac{n-1}{L-1}(\tau_{\mathrm{max}}-\tau_{\mathrm{min}})$, with $\tau_{\mathrm{min}}$ and $\tau_{\mathrm{max}}$ denoting the minimum and maximum times, and $L$ is the number of steps. Let us denote the uncertainty of the estimated energy as $\delta\omega$. According to the time–energy uncertainty relation, $1/\Delta\tau\simeq\delta\omega$, where $\Delta\tau=(\tau_{\mathrm{max}}-\tau_{\mathrm{min}})/(L-1)$ is the time interval, which can be chosen to match the desired estimation precision.
The peak of $f(\omega)$ appears at $\omega$ corresponding to the energy difference $E^{\mathrm{g}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}}$. Since $E^{n}_{q,\mathrm{P}}$ is known, the ground-state energy in the subspace corresponding to eigenvalue $q^{\prime}$ can be determined. From this procedure, we successfully estimate the ground-state energy in the subspace with $q^{\prime}$. To estimate the ground-state energy in another subspace with $q^{\prime\prime}$, we could perform the above procedure, say $q^{\prime\prime}$. Thanks to the second assumption of our protocol, we can perform this search across all subspaces within polynomial time. Similarly, the first-excited-state energy in the subspace of $q^{\prime}$ can be estimated by preparing a superposition of the first excited state and the reference state of $\hat{H}_{\mathrm{D}}$ in the subspace of eigenvalues $q^{\prime}$ and $q$ respectively, as the initial state.
@@
Furthermore, if non-adiabatic transitions occur during the ASP, peaks appear in $f(\omega)$ at frequencies corresponding to the energy difference between the excited state and the reference state of $\hat{H}_{\mathrm{P}}$, such as $E^{\mathrm{e_{1}}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}},E^{\mathrm{e_{2}}}_{q^{\prime},\mathrm{P}}-E^{n}_{q,\mathrm{P}},\dots$, where $E^{\mathrm{e}_{i}}_{q^{\prime},\mathrm{P}}$ denotes the $i$-th excited-state energy in the subspace associated with eigenvalue $q^{\prime}$. These excitation energies can also be estimated since the trivial eigenvalue is known in advance. This implies that our method does not require careful tuning of the ASP execution time $T$ in advance. This is in stark contrast to the conventional scheme, where the ground state is prepared adiabatically and the ground-state energy is estimated by measuring the expectation value of the Hamiltonian. In such conventional methods, even a small non-adiabatic transition can significantly reduce the fidelity of the prepared state with respect to the true ground state of the problem Hamiltonian, resulting in inaccurate energy estimation.
Specifically, the problem Hamiltonian can be decomposed as $\hat{H}_{\mathrm{P}}=\sum_{j}c_{j}\hat{P}_{j}$, where $c_{j}$ is a real number and $\hat{P}_{j}$ is a tensor product of the Pauli operators. In conventional methods, the ground-state energy is estimated by preparing the ground state via ASP and then measuring the expectation value $\langle\hat{H}_{\mathrm{P}}\rangle=\sum_{j}c_{j}\langle\hat{P}_{j}\rangle$. To ensure high fidelity with the true ground state, the total execution time $T_{\rm conv}$ must satisfy the adiabatic condition [36, 37, 38, 39]. The time-dependent Hamiltonian in the conventional ASP method is typically given as $\hat{H}_{\rm conv}=(1-s)\hat{H}_{\mathrm{D}}+s\hat{H}_{\mathrm{P}}$, where $s=t/T_{\rm conv}$ the normalized time, $t$ is the physical time, and $T_{\rm conv}$ is the total execution time of ASP. The adiabatic condition is then explained as
$\max_{s\in[0,1]}\frac{|\bra{\psi^{\mathrm{e}_{n}}(s)}\dot{\hat{H}}(s)\ket{\psi^{\mathrm{g}}(s)}|}{{|E^{\mathrm{e}_{n}}(s)-E^{\mathrm{g}}(s)|}^{2}}\ll T_{\rm conv},$ (12)
where |$\ket{\psi^{\mathrm{e}_{n}}(s)}$ and |$\ket{\psi^{\rm g}(s)}$ denote the $n$-th excited and ground state of the instantaneous Hamiltonian $\hat{H}(s)$, respectively, and $E^{\mathrm{e}_{n}}(s)$ and $E^{\rm g}(s)$ are their corresponding eigenenergies. $\dot{\hat{H}}(s)$ is the derivative of $\hat{H}(s)$ with respect to the normalized time $s$. In practice, neither the transition matrix elements nor the energy gaps of the Hamiltonian are known in advance, making it difficult to determine a suitable value of $T_{\rm conv}$ that ensures a high fidelity beforehand. In contrast, our method allows for accurate energy estimation even in the presence of non-adiabatic transitions, by utilizing interference between the ground state and a reference state. Therefore, our approach eliminates the need for fine-tuning of the ASP runtime and provides a practical advantage over the conventional expectation-value-based method. It should be noted, however, if non-adiabatic transitions significantly reduce the ground state population, both our method and the conventional quantum phase estimation fail to determine the ground-state energy.
Regarding the reference state, since it is chosen from a subspace of size $O(\mathrm{poly}(N))$, non-adiabatic transitions can be suppressed byu setting $T$ appropriately within polynomial time. Therefore, non-adiabatic transitions involving the reference state are considered negligible in our analysis.
@@
III Result
We estimate the ground-state energy of the spin1/212-1/2- 1 / 2 Heisenberg model using our proposed method through numerical simulations. In this study, we consider a 4-qubit system. The problem Hamiltonian is defined as the Heisenberg model with an additional inhomogeneous longitudinal magnetic field to lift the degeneracy: